OpenStudy (anonymous):

(g + 4) - 3g = 1 + g SHOW WORK PLEASE

OpenStudy (anonymous):

\[(g+4)-3g=1+g\]add 3g to both sides:\[(g+4)=1+4g\]remove the parenthesis and subtract the g from both sides:\[4=1+3g\]subtract 1 from both sides:\[3=3g\]divide both sides by 3 to isolate g:\[1=g\] \[\large g=1~~\huge\ddot\smile\]

OpenStudy (anonymous):

@claudiavv hope this helps :)

OpenStudy (anonymous):

you can also check your work by putting in 1 for g to see if it works: \[~~~~(g+4)-3g=1+g\]\[(1+4)-3\times1=1+1\]\[~~~~~~~~~~~~~~~5-3=2\]\[~~~~~~~~~~~~~~~~~~~~~2=2~~\checkmark\]

OpenStudy (anonymous):

yummydum did it right!

OpenStudy (anonymous):

go yummydum!!!!!!!!!

OpenStudy (anonymous):

lol...thanks Cheetah :)

OpenStudy (anonymous):

LIKE THE FACE YOU HAVE THERE!